# On the use of galaxy catalogs in gravitational-wave parameter estimation

### v1

@@ 3.4 Performance on simulated data
In addition to the real GW events above, we also tested our method on simulated BNS mergers, where we found mixed results. Our simulations consisted of 250 BNS mergers sampled with parameters as described in Table 3 using the IMRPhenomD waveform model (Husa et al., 2016; Khan et al., 2016; Pürrer et al., 2023; Morisaki et al., 2023) and a network of LIGO Hanford, LIGO Livingston, and Virgo at O4 design sensitivity [^4] We use the aligo_O4high.txt and avirgo_O4high_NEW.txt noise curves from Abbott et al. (2020a)., simulated in Gaussian noise. We place the simulated mergers at the locations of galaxies in the catalog. We further enforce a network optimal signal-to-noise ratio (S/N) cut of 8, to avoid very poorly localized events. 131 events pass this S/N cut, of which ten events were removed due to problems with sky map generation from posterior samples (for either the 2D or 3D localizations), resulting in 121 simulated BNS mergers which we analyze below. [^5] These ten events did not show obvious trends in their properties, but we acknowledge that this could result in a $\approx 8$% bias in our cumulative results. These simulations are intended to represent a range of realistic BNS detections for which our method is relevant.
@@
However, we found that the crossmatch of the galaxy catalog with the initial parameter estimation localization (ie., Step 2 described in Sec. 2.2), when sorted by the volume searched probability, performs similarly to our method using a 3D catalog. Since this crossmatch is performed with the sky map, which is generated from a kernel density estimate using a subset of the posterior samples, the resulting probabilities at each galaxy location are, in principle, not as precise estimates of the likelihood compared to our method of sampling directly at the discrete position of each galaxy. In practice, our results show that this discrete sampling does not result in significant improvement. Indeed, for about 13% of our simulated events, the true host was not sampled by Bilby when localized with the 3D catalog. This was most prevalent for simulated events at true distances larger than 120 Mpc, where the number of galaxies in the 3D catalog would often be over $10^{4}$.
@@ 4 Discussion and conclusion
The completeness of the input galaxy catalog is one of the most important factors in using this method. Existing catalogs such as NED-LVS begin to fall below 70% K-band completeness at around 350 Mpc ($z\sim 0.075$) (Cook et al., 2023). Current and future surveys such as SDSS-V (York et al., 2000; Kollmeier et al., 2017), DESI (DESI Collaboration et al., 2016), LSST at the Vera Rubin Observatory (LSST Science Collaboration et al., 2009; Ivezić et al., 2019), Euclid (Laureijs et al., 2011), SPHEREx (Doré et al., 2014), and Roman (Spergel et al., 2015) will push the forefront of survey science, ensuring that galaxy catalogs will continue to deepen. Future catalogs may enable the use of our method for mergers at higher redshifts. In a similar vein, powerful upcoming follow-up instruments such as the Vera Rubin Observatory and DSA-2000 (Hallinan et al., 2019) might be able to produce on-the-fly galaxy catalogs for a given initial localization which may be deeper than existing catalogs (Morsony et al., 2024), enabling an iterative process between GW and EM observers. Also, if this method can be sped up to $\mathcal{O}$(minute) latencies with further GW parameter estimation improvements, rapid catalog-informed localizations could be useful for compute- and storage-limited radio telescope follow-up (e.g., the Long Wavelength Array (Anderson et al., 2018)), where beamforming on the sky is currently an expensive bottleneck.
@@
We would like to acknowledge useful discussions with Gregg Hallinan, Dave Cook, Salvatore Vitale, and Leo Singer. G. M. acknowledges the support of the National Science Foundation and the LIGO Laboratory. C.-J, H. acknowledges the support from NASA Grant 80NSSC23M0104 and the Nevada Center for Astrophysics. LIGO was constructed by the California Institute of Technology and Massachusetts Institute of Technology with funding from the National Science Foundation and operates under cooperative agreement PHY-0757058. The authors are grateful for computational resources provided by the LIGO Lab and supported by NSF Grants PHY-0757058 and PHY-0823459.

### v2

@@ 3.4 Performance on simulated data
In addition to the real GW events above, we also tested our method on simulated BNS mergers, where we found mixed results. Our simulations consisted of 250 BNS mergers sampled with parameters as described in Table 3 using the IMRPhenomD waveform model (Husa et al., 2016; Khan et al., 2016; Pürrer et al., 2023; Morisaki et al., 2023) and a network of LIGO Hanford, LIGO Livingston, and Virgo at O4 design sensitivity [^4] We use the aligo_O4high.txt and avirgo_O4high_NEW.txt noise curves from Abbott et al. (2020a)., simulated in Gaussian noise. We place the simulated mergers at the locations of galaxies in the catalog, enforcing only that a host’s Ks-band luminosity is greater than 10% of the Milky Way luminosity to avoid potentially clustered low-mass galaxies [^5] This reduces the catalog size from 1.9 million to 1.4 million galaxies, a relatively minor effect.. We further enforce a network optimal signal-to-noise ratio (S/N) cut of 8, to avoid very poorly localized events. 131 events pass this S/N cut, of which ten events were removed due to problems with sky map generation from posterior samples (for either the 2D or 3D localizations), resulting in 121 simulated BNS mergers which we analyze below. [^6] These ten events did not show obvious trends in their properties, but we acknowledge that this could result in a $\approx 8$% bias in our cumulative results. These simulations are intended to represent a range of realistic BNS detections for which our method is relevant.
@@
However, we found that the crossmatch of the galaxy catalog with the initial parameter estimation localization (ie., Step 2 described in Sec. 2.2), when sorted by the volume searched probability, performs similarly to our method using a 3D catalog [^7] We do not perform a mass or luminosity weighting in using the 3D catalog, only enforcing that hosts have $L_{\rm Ks}>0.1L_{\rm MW}$ to match the simulated population.. Since this crossmatch is performed with the sky map, which is generated from a kernel density estimate using a subset of the posterior samples, the resulting probabilities at each galaxy location are, in principle, not as precise estimates of the likelihood compared to our method of sampling directly at the discrete position of each galaxy [^8] See e.g., Magaña Hernandez et al. (2024) for a demonstration of the possible variance in distance probability density when estimated from posterior samples using different methods. While this likely has a small effect on our results, it is important to keep in mind for counterpart searches using catalog cross-matches from density-estimated posteriors.. In practice, our results show that this discrete sampling does not result in significant improvement. Indeed, for about 13% of our simulated events, the true host was not sampled by Bilby when localized with the 3D catalog. This was most prevalent for simulated events at true distances larger than 120 Mpc, where the number of galaxies in the 3D catalog would often be over $10^{4}$.
@@ 4 Discussion and conclusion
The completeness of the input galaxy catalog is one of the most important factors in using this method. Existing catalogs such as NED-LVS begin to fall below 70% K-band completeness at around 350 Mpc ($z\sim 0.075$) (Cook et al., 2023). Current and future surveys such as SDSS-V (York et al., 2000; Kollmeier et al., 2017), DESI (DESI Collaboration et al., 2016), LSST at the Vera Rubin Observatory (LSST Science Collaboration et al., 2009; Ivezić et al., 2019), Euclid (Laureijs et al., 2011), SPHEREx (Doré et al., 2014), and Roman (Spergel et al., 2015) will push the forefront of survey science, ensuring that galaxy catalogs will continue to deepen. Future catalogs may enable the use of our method for mergers at higher redshifts. In a similar vein, powerful upcoming follow-up instruments such as the Vera Rubin Observatory and DSA-2000 (Hallinan et al., 2019) might be able to produce on-the-fly galaxy catalogs for a given initial localization which may be deeper than existing catalogs (Morsony et al., 2024), enabling an iterative process between GW and EM observers. Also, if this method can be sped up to $\mathcal{O}$(minute) latencies with further GW parameter estimation improvements, rapid catalog-informed localizations could be useful for compute- and storage-limited radio telescope follow-up (e.g., the Long Wavelength Array; Anderson et al. 2018), where beamforming on the sky is currently an expensive bottleneck.
@@
We would like to acknowledge useful discussions with Gregg Hallinan, Dave Cook, Sylvia Biscoveanu, Salvatore Vitale, and Leo Singer. G. M. acknowledges the support of the National Science Foundation and the LIGO Laboratory. C.-J, H. acknowledges the support from NASA Grant 80NSSC23M0104 and the Nevada Center for Astrophysics. LIGO was constructed by the California Institute of Technology and Massachusetts Institute of Technology with funding from the National Science Foundation and operates under cooperative agreement PHY-0757058. The authors are grateful for computational resources provided by the LIGO Lab and supported by NSF Grants PHY-0757058 and PHY-0823459.